OpenStudy (anonymous):

How do you Rewrite an expression without using grouping symbols ? 11(23 – x) A.)11 • 23 – x B.) 1123 – 11x C.)11 • 23 – 11x D.) 11 + 23 – 11 + x

OpenStudy (slaaibak):

Example?

OpenStudy (anonymous):

11(23 – x)

OpenStudy (hba):

–6(8 + 12) =(-6 )*( 8) + (-6) * (12)

OpenStudy (slaaibak):

Use the distributive law: a * (b + c) = (a * b) + (a * c)